OpenStudy (anonymous):

The ratio of similarity of two house lots is 3:1.8. If the smaller lot is 93 yards around, what is the perimeter of the larger lot? 115 125 155 279

OpenStudy (anonymous):

3:1.8 is 5:2 but i dont know what to do now?

OpenStudy (anonymous):

Solve the following for L, the larger lot's perimeter.\[\frac{3}{\frac{18}{10}}=\frac{L}{93} \]\[L=155 \]

OpenStudy (anonymous):

oh i getn it now i really appreciate it now i can work these problems mysef, thanks for great teaching!

OpenStudy (anonymous):

Thank you for the medal, and you are welcome.